my problem is this.
Solve the problem.The sum of the measures of the angles in any triangle is 180 degrees. In triangle ABC, angles A and B have the same measure, while angle C is 42 degrees larger than each of the other two angles. Find the measure of angle C.
i don't understand how to find the measure of angle C.

Answer by Alan3354(69443)

In triangle ABC, angles A and B have the same measure, while angle C is 42 degrees larger than each of the other two angles. Find the measure of angle C.
================
A = B
C = A + 42
----
A + B + C = 180
A + A + A+42 = 180
3A + 42 = 180
3A = 138
A = B = 46 degs
C = A + 42
C = 88 degs